How many grams of solid sodium hydroxide is required to make 500.00 mL of a 0.79 M NaOH solution?

Answer with two decimal places.

What is the concentration of a sodium hydroxide solution if 0.2723 g of potassium hydrogen phthalate (KHP- MW 204.22 g/mol) requires 10.02 mL of NaOH
solution to reach the phenolphthalein endpoint?

Answer with three decimal places.

How many milliliters of 0.1200 M NaOH is needed to titrate 0.2100 g KHP (MW 204.22 g/mol) to the phenolphthalein end
point?

Answer with two decimal places.
Answer #1

1)

Molarity = moles / volume

0.79 = moles / 0.500

moles = 0.395

moles = mass / molar mass

0.395 = mass / 40

mass of sodium hydroxide = 15.80 g

2)

Answer with three decimal places.

mass of KHP = 0.2723 g

moles of KHP = 0.2723 / 204.22 = 1.33 x 10^-3

moles of KHP = moles of NaOH

1.33 x 10^-3 = M x 10.02 x 10^-3

M = 0.133

concentration of NaOH = 0.133 M

3)

moles of KHP = 0.2100 / 204.22 = 1.03 x 10^-3 mol

1.03 x 10^-3 = 0.12 x V

V = 0.00857 L

volume of NaOH = 8.57 mL